OpenStudy (anonymous):

Solve by factoring. Mark the solutions for: 2x2 + 3x - 14 = 0 A. 2, -2 B. -2, 2 C. 2, -7/2 D. -2, 7/2

OpenStudy (callisto):

2x^2 + 3x - 14 = 0 (x-2) (2x+7) =0 Put x-2 =0 or 2x+7 =0 x=2 or x=-7/2 I would appreciate you can ask what you don't understand and show us your workings